# Data-Efficient Time-Dependent PDE Surrogates: Graph Neural Simulators vs Neural Operators

### v1

@@ Abstract
Neural operators (NOs) approximate mappings between infinite-dimensional function spaces; however, they generally demand large datasets and can struggle to generalize when training data are scarce. Moreover, many NO formulations do not explicitly encode the causal, local-in-time structure of physical evolution. Autoregressive models that predict the next time-step inherently preserve causality, yet they are prone to rapid error accumulation. In this work, we employ Graph Neural Simulators (GNS) - a general framework in message-passing graph neural networks - coupled with explicit numerical time-stepping schemes to construct accurate and efficient forward models that learn PDE solutions by directly modeling the instantaneous time derivative. We evaluate the efficacy and specifically the generalization capabilities of our framework on three canonical PDE systems: (1) the 2D Burgers’ scalar equation, (2) the 2D coupled Burgers’ vector equation, and (3) the 2D Allen–Cahn equation. Rigorous evaluations across all cases demonstrate that GNS significantly improves data efficiency, achieving higher generalization accuracy with substantially fewer training trajectories compared to neural operator baselines such as DeepONet and FNO. Notably, GNS achieves consistently under 1% relative $L_{2}$ errors with only 3% of the available trajectories (30 training samples out of 1000) across all three PDE systems. Moreover, it substantially reduces error accumulation over extended temporal horizons compared to its neural operator autoregressive (AR) baselines: averaged across the three PDE systems, GNS reduces AR error by 82.48% relative to FNO AR and by 99.86% relative to DON AR. We further introduce a PCA+KMeans trajectory selection strategy that enhances performance in the low-data setting. Our results indicate that combining graph-based local inductive biases with conventional time integrators yields accurate, physically consistent, and scalable surrogate models for time-dependent PDEs. Code and data are available at the project repository: https://github.com/Centrum-IntelliPhysics/GNS_vs_NOs.git.
@@ 1 Introduction
Figure 1: Comparison of neural network architectures for neural operators in scientific applications. Fully connected networks (FNNs) flatten mesh data, discarding spatial structure and compressing high-dimensional inputs into narrow hidden layers, which limits expressivity. Convolutional neural networks (CNNs) require interpolation to uniform grids at both input and output, sacrificing fidelity on unstructured (adaptive) meshes and introducing errors tied to interpolation order. Graph neural networks (GNNs), by contrast, natively represent mesh nodes and edges as graphs, preserving geometry, topology, and physical interactions. Their intrinsic alignment with mesh discretizations, ability to exploit locality, and data efficiency make GNNs the natural choice for Scientific Machine Learning, where irregular geometries and fine discretizations are the norm.
@@
Building on the flexibility of GNNs, Graph Neural Simulators (GNS) have been proposed as powerful data-driven surrogates for time-dependent physical systems [28, 29]. By leveraging message passing [30] to encode local temporal updates, GNS models excel in capturing the evolution of PDE-governed systems on irregular domains, particularly in settings with unstructured meshes, particle-based discretizations, or evolving geometries. This locality-driven formulation enables improved generalization across domains, a challenge where NOs often struggle. Unlike NOs that directly learn global solution operators across function spaces, GNS emphasizes local dynamics by propagating information through graph neighborhoods, offering a more physically intuitive representation. However, this approach relies on autoregressive rollout for long-term predictions, which can introduce error accumulation over extended time horizons. Thus, neural operators and GNN-based simulators represent two complementary paradigms — global operator learning versus local dynamics learning, each with distinct strengths and limitations for modeling complex physical phenomena.
Building on recent work that introduced a DeepONet-based surrogate model for accurate temporal extrapolation in time-dependent PDEs [31], we employ a GNS framework that learns to approximate the instantaneous time derivative of autonomous PDE systems. By reformulating the temporal evolution problem as learning the right-hand side function, our approach leverages the inherent spatial locality of GNS architectures to model derivative terms while enabling stable long-term integration through established numerical time-stepping schemes. This methodology combines the geometric flexibility of graph-based representations with the temporal stability of classical numerical methods. In this regard, we make the following key contributions:
@@
  • We introduce a systematic training sample selection strategy (specifically for low data regime) that combines principal component analysis (PCA) [32] with KMeans clustering [33] to ensure sufficient representation of the full dataset through the selected traning samples, demonstrating improved performance over pure random sampling approaches.
@@ 3 Results
To ensure a fair and comprehensive evaluation, we benchmark both GNS and NO frameworks on three canonical PDE systems that span a diverse range of complex, nonlinear, high-dimensional, and spatiotemporal dynamics: (1) the 2D Burgers’ equation with a scalar field solution, (2) the 2D coupled Burgers’ equation with a vector field solution, and (3) the 2D Allen-Cahn equation. For the comparison of inference errors, we consider a relative $L_{2}$ error metric, defined as
@@
Figure 2 depicts the temporal error accumulation observed across all frameworks and examples considered in this study. Across all three PDE systems, GNS (black) consistently exhibits the lowest error growth with a noticeable plateauing of errors, indicating that the GNS surrogate demonstrates accurate, reliable, and stable performance over extended temporal horizons. This is followed by FNO-FR (blue) and FNO-AR (orange), with FNO-FR performing better than FNO-AR for longer timesteps except in the 2D Burgers’ scalar case, where FNO-AR surpasses FNO-FR beyond $t=0.7$. Note that FNO-FR incurs appreciable errors for the initial few timesteps across all three cases, beyond which it outperforms FNO-AR. Finally, DeepONet-based frameworks incur substantially higher errors than their counterparts, i.e., FNO and GNS. The primary reason is that DeepONet generally requires significantly more training data than the other methods to construct an accurate set of basis functions and coefficients. Among the DeepONet frameworks, the time-integrator variants, TI(L)-DON (magenta) and TI-DON (green), perform better than DON-FR (red) for the initial timesteps until $t=0.45$ or $t=0.5$. Beyond this point, DON-FR surpasses these frameworks. The aforementioned observations translate directly to the comparison plots in subsequent sections where we discuss individual cases separately and examine this phenomenon through both qualitative and quantitative assessment of generalization performance via relative $L_{2}$ test errors during inference. Details regarding the computational costs incurred during training and inference are provided in Table A1 in the Supplementary Information (SI). Additional information on the network architectures, as well as further results on predicted solution contours, is provided in SI B and SI C, respectively. The code to reproduce the experiments is publicly available at https://github.com/Centrum-IntelliPhysics/GNS_vs_NOs.git.
Table 1: Ground truth dataset generation details for the three PDE cases. The data generation script for both 2D Burgers’ PDEs was adapted from [36]. The generalization error of all frameworks for each PDE is evaluated using training data sizes of 30, 50, 70, and 100 samples. Training trajectories are selected using a combined PCA and KMeans approach.
@@
Figure 2: Temporal evolution of error accumulation across different frameworks for all examples, trained with 30 trajectories. Left: comparison of all NO baselines with the GNS surrogate model. Right: comparison of the GNS surrogate model with the FNO variants (FNO-FR and FNO-AR). In the 2D coupled Burgers’ case, figures (c) and (d) correspond to the $u$-velocity field, while figures (e) and (f) correspond to the $v$-velocity field.
@@ 3.1 Two-Dimensional Burgers’ Equation with Scalar Output Field
Figure 3: 2D Burgers’ equation with a scalar output field: Representative sample from the 1000 testing trajectories demonstrating the performance of all ML frameworks trained on 30 trajectories.
@@ 3.2 Two-Dimensional Coupled Burgers’ Equation with Vector Output Field
The second example we consider is the two-dimensional coupled Burgers’ equation with vector-valued outputs, representing the velocity components $u(x,y,t)$ and $v(x,y,t)$. This system can be regarded as a simplification of the incompressible Navier–Stokes equations that does not account for pressure effects, extending the scalar Burgers’ equation by introducing two interacting velocity fields. In doing so, it captures richer transport phenomena by balancing nonlinear cross-advection with viscous diffusion. The governing equations are:
@@
Figure 4: 2D coupled Burgers’ equation with a vector output field: Performance of all frameworks on 30 training trajectories, evaluated on 1000 trajectories for predicting the $u$-velocity field of a representative sample.
Figure 5: 2D coupled Burgers’ equation with a vector field: Performance of all frameworks on 30 training trajectories, evaluated on 1000 trajectories for predicting the $v$-velocity field of a representative sample.
In a similar spirit, we first present a qualitative comparison of the prediction test errors incurred during inference on 1000 trajectories, trained on 30 trajectories. The results are shown in Figs. 4 and 5 for the $u$ and $v$ velocity fields, respectively. While the contours are displayed separately for the two fields, the overall trends remain consistent across both. As before, DON-AR exhibits errors an order of magnitude larger than all other methods, which can be attributed to its high sensitivity to error accumulation, exacerbated by the limited availability of training data. Among the other DeepONet-based variants, the TI-based frameworks (TI(L)-DON and TI-DON) outperform DON-FR during the initial rollout (up to $t=0.4$). Beyond this temporal horizon, DON-FR surpasses the TI-based methods and achieves lower test errors. Within the TI family, TI(L)-DON consistently outperforms the vanilla TI-DON due to its ability to adapt to local solution features, an advantage that becomes especially useful when training data is scarce.
@@
Table 3 summarizes the overall relative $L_{2}$ inference errors for predicting the $u$ and $v$ velocity fields using 30, 50, 70, and 100 training trajectories. The observed trends align closely with the qualitative comparisons in Figs. 4 and 5. As expected, DON-AR performs the worst, with relative $L_{2}$ errors exceeding $100\%$ across all cases, indicating that its predicted solutions deviate entirely from the ground truth. Among the remaining DeepONet-based frameworks, the TI variants (TI-DON and TI(L)-DON) achieve slightly lower errors than DON-FR when trained with only 30 trajectories and perform comparably for 50 and 70 trajectories. However, when sufficient training data is provided, DON-FR yields more accurate basis function and coefficient representations of the output solution space, ultimately surpassing the TI variants in predictive accuracy.
@@ 3.3 Two-dimensional Allen–Cahn Equation
Figure 6: 2D Allen-Cahn Equation: Performance of all frameworks on 30 training trajectories, evaluated on 1000 trajectories for a representative sample.

### v2

# Data-Efficient Time-Dependent PDE Surrogates: Graph Neural Simulators vs. Neural Operators

@@ Abstract
Developing accurate, data-efficient surrogate models is central to advancing AI for Science. Neural operators (NOs), which approximate mappings between infinite-dimensional function spaces using conventional neural architectures, have gained popularity as surrogates for systems driven by partial differential equations (PDEs). However, their reliance on large datasets and limited ability to generalize in low-data regimes hinder their practical utility. We argue that these limitations arise from their global processing of data, which fails to exploit the local, discretized structure of physical systems. To address this, we propose Graph Neural Simulators (GNS) as a principled surrogate modeling paradigm for time-dependent PDEs. GNS leverages message-passing combined with numerical time-stepping schemes to learn PDE dynamics by modeling the instantaneous time derivatives. This design mimics traditional numerical solvers, enabling stable long-horizon rollouts and strong inductive biases that enhance generalization. We rigorously evaluate GNS on four canonical PDE systems: (1) 2D scalar Burgers’, (2) 2D coupled Burgers’, (3) 2D Allen–Cahn, and (4) 2D nonlinear shallow-water equations, comparing against state-of-the-art NOs including Deep Operator Network (DeepONet) and Fourier Neural Operator (FNO). Results demonstrate that GNS is markedly more data-efficient, achieving $1\%$ relative $L_{2}$ error using only 3% of available trajectories, and exhibits dramatically reduced error accumulation over time (82.5% lower autoregressive error than FNO, 99.9% lower than DeepONet). To choose the training data, we introduce a PCA+KMeans trajectory selection strategy. These findings provide compelling evidence that GNS, with its graph-based locality and solver-inspired design, is the most suitable and scalable surrogate modeling framework for AI-driven scientific discovery. Code and data are available at the project repository: https://github.com/Centrum-IntelliPhysics/GNS_vs_NOs.git.
@@ 1 Introduction
Figure 1: Comparison of neural network architectures for neural operators in scientific applications. Fully connected neural networks (FNNs) flatten mesh data, discarding spatial structure and compressing high-dimensional inputs into narrow hidden layers, which limits expressivity. Convolutional neural networks (CNNs) require interpolation to uniform grids at both input and output, sacrificing fidelity on unstructured (adaptive) meshes and introducing errors tied to interpolation order. Graph neural networks (GNNs), by contrast, natively represent mesh nodes and edges as graphs, preserving geometry, topology, and physical interactions. Their intrinsic alignment with mesh discretizations, ability to exploit locality, and data efficiency make GNNs the natural choice for Scientific Machine Learning, where irregular geometries and fine discretizations are the norm.
@@
Figure 2: A schematic of the Graph Neural Simulator (GNS) architecture (adapted from [28]).
Building on the flexibility of GNNs, Graph Neural Simulators (GNS) have been proposed as powerful data-driven surrogates for time-dependent physical systems [28, 29]. For completeness, a schematic of the GNS architecture, adapted from [28], is presented in Fig. 2. By leveraging message passing [30] to encode local temporal updates, GNS models excel in capturing the evolution of PDE-governed systems on irregular domains, particularly in settings with unstructured meshes, particle-based discretizations, or evolving geometries. This locality-driven formulation enables improved generalization across domains, a challenge where NOs often struggle. Unlike NOs that directly learn global solution operators across function spaces, GNS emphasizes local dynamics by propagating information through graph neighborhoods, offering a more physically intuitive representation. Thus, neural operators and GNN-based simulators represent two complementary paradigms — global operator learning versus local dynamics learning, each with distinct strengths and limitations for modeling complex physical phenomena.
Building on recent work that introduced a DeepONet-based surrogate model for accurate temporal extrapolation in time-dependent PDEs [31], we employ a GNS framework that learns to approximate the instantaneous time derivative of autonomous PDE systems (see Fig. 2). By reformulating the temporal evolution problem as learning the right-hand side (RHS) function, our approach leverages the inherent spatial locality of GNS architectures to model derivative terms while enabling stable long-term integration through established numerical time-stepping schemes. This methodology combines the geometric flexibility of graph-based representations with the temporal stability of classical numerical methods. In this regard, we make the following key contributions:
@@
  • We introduce a systematic training sample selection strategy (specifically for low data regime) that combines principal component analysis (PCA) [32] with KMeans clustering [33] to ensure sufficient representation of the full dataset through the selected training samples, demonstrating improved performance over pure random sampling approaches.
@@ 3 Results
To ensure a fair and comprehensive evaluation, we benchmark both GNS and NO frameworks on four canonical PDE systems that cover a diverse range of complex, nonlinear, high-dimensional, and spatiotemporal dynamics: (1) the 2D Burgers’ equation with a scalar field solution, (2) the 2D coupled Burgers’ equation with a vector field solution, (3) the 2D Allen–Cahn equation, and (4) the 2D nonlinear shallow water equations (SWE). Since we are comparatively assessing the generalization performance of the GNS surrogate against other NO baselines for time-dependent PDEs, we restrict our study to uniform grids, as most NO architectures (e.g., FNO) require a uniform grid domain. For the comparison of inference errors, we consider a relative $L_{2}$ error metric, defined as
@@
Figure 3 depicts the temporal error accumulation observed across all frameworks and the first three PDE examples considered in this study. Across these PDE systems, GNS (black) consistently exhibits the lowest error growth with a noticeable plateauing of errors, indicating that the GNS surrogate demonstrates accurate, reliable, and stable performance over extended temporal horizons. This is followed by FNO-FR (blue) and FNO-AR (orange), with FNO-FR performing better than FNO-AR for longer timesteps except in the 2D Burgers’ scalar case, where FNO-AR surpasses FNO-FR beyond $t=0.7$. Note that FNO-FR incurs appreciable errors for the initial few timesteps across all three cases, beyond which it outperforms FNO-AR. Finally, DeepONet-based frameworks incur substantially higher errors than their counterparts, i.e., FNO and GNS. The primary reason is that DeepONet generally requires significantly more training data than the other methods to construct an accurate set of basis functions and coefficients. Among the DeepONet frameworks, the time-integrator variants, TI(L)-DON (magenta) and TI-DON (green), perform better than DON-FR (red) for the initial timesteps until $t=0.45$ or $t=0.5$. Beyond this point, DON-FR surpasses these frameworks. The aforementioned observations translate directly to the comparison plots in subsequent sections where we discuss individual cases separately and examine this phenomenon through both qualitative and quantitative assessment of generalization performance via relative $L_{2}$ test errors during inference. Details regarding the computational costs incurred during training and inference are provided in Table A1 in the Supplementary Information (SI). Additional information on the network architectures, as well as further results on predicted solution contours, is provided in SI B and SI C, respectively. The code to reproduce the experiments is publicly available at https://github.com/Centrum-IntelliPhysics/GNS_vs_NOs.git.
Table 1: Ground truth dataset generation details for the four PDE cases. The data generation script for the 2D Burgers’ examples and the 2D nonlinear SWE example was adapted from [36]. The generalization error of all frameworks for each PDE is evaluated using training data sizes of 30, 50, 70, and 100 samples. Training trajectories are selected using a combined PCA and KMeans approach.
@@
2D nonlinear
SWE
Matérn Periodic 0.1 0.2 500 30 500 $64\times 64$ 101
50
70
100
Figure 3: Temporal evolution of error accumulation across different frameworks for the 2D Burgers’ scalar, 2D Burgers’ vector, and 2D Allen-Cahn examples, trained with 30 trajectories. Left: comparison of all NO baselines with the GNS surrogate model. Right: comparison of the GNS surrogate model with the FNO variants (FNO-FR and FNO-AR). In the 2D coupled Burgers’ case, figures (c) and (d) correspond to the $u$-velocity field, while figures (e) and (f) correspond to the $v$-velocity field.
@@ 3.1 Two-Dimensional Burgers’ Equation with Scalar Output Field
Figure 4: 2D Burgers’ equation with scalar output field. Performance of all frameworks trained on 30 trajectories and evaluated on 1000. Top row: ground truth solution contours; subsequent rows: absolute error contours for each framework.
@@ 3.2 Two-Dimensional Coupled Burgers’ Equation with Vector Output Field
The second example we consider is the two-dimensional coupled Burgers’ equation with vector-valued outputs, representing the velocity components $u(x,y,t)$ and $v(x,y,t)$. This system can be regarded as a simplification of the incompressible Navier–Stokes equations that does not account for pressure effects, extending the scalar Burgers’ equation by introducing two interacting velocity fields. In doing so, it captures richer transport phenomena by balancing nonlinear cross-advection with viscous diffusion. The system is defined on the spatial domain $(x,y)\in[0,1]^{2}$ over the time interval $t\in[0,1]$, and its governing equations are:
@@
Figure 5: 2D coupled Burgers’ equation with a vector output field: Performance of all frameworks on 30 training trajectories, evaluated on 500 trajectories for predicting the $u$-velocity field of a representative sample. Top row: ground truth solution contours; subsequent rows: absolute error contours for each framework.
Figure 6: 2D coupled Burgers’ equation with a vector field: Performance of all frameworks on 30 training trajectories, evaluated on 500 trajectories for predicting the $v$-velocity field of a representative sample. Top row: ground truth solution contours; subsequent rows: absolute error contours for each framework.
In a similar spirit, we first present a qualitative comparison of the prediction test errors incurred during inference on 1000 trajectories, trained on 30 trajectories. The results are shown in Figs. 5 and 6 for the $u$ and $v$ velocity fields, respectively. While the contours are displayed separately for the two fields, the overall trends remain consistent across both. As before, DON-AR exhibits errors an order of magnitude larger than all other methods, which can be attributed to its high sensitivity to error accumulation, exacerbated by the limited availability of training data. Among the other DeepONet-based variants, the TI-based frameworks (TI(L)-DON and TI-DON) outperform DON-FR during the initial rollout (up to $t=0.4$). Beyond this temporal horizon, DON-FR surpasses the TI-based methods and achieves lower test errors. Within the TI family, TI(L)-DON consistently outperforms the vanilla TI-DON due to its ability to adapt to local solution features, an advantage that becomes especially useful when training data is scarce.
@@
Table 3 summarizes the overall relative $L_{2}$ inference errors for predicting the $u$ and $v$ velocity fields using 30, 50, 70, and 100 training trajectories. The observed trends align closely with the qualitative comparisons in Figs. 5 and 6. As expected, DON-AR performs the worst, with relative $L_{2}$ errors exceeding $100\%$ across all cases, indicating that its predicted solutions deviate entirely from the ground truth. Among the remaining DeepONet-based frameworks, the TI variants (TI-DON and TI(L)-DON) achieve slightly lower errors than DON-FR when trained with only 30 trajectories and perform comparably for 50 and 70 trajectories. However, when sufficient training data is provided, DON-FR yields more accurate basis function and coefficient representations of the output solution space, ultimately surpassing the TI variants in predictive accuracy.
@@ 3.3 Two-Dimensional Allen–Cahn Equation
Figure 7: 2D Allen-Cahn Equation: Performance of all frameworks on 30 training trajectories, evaluated on 1000 trajectories for a representative sample. Top row: ground truth solution contours; subsequent rows: absolute error contours for each framework.
@@
3.4 Two-Dimensional nonlinear Shallow Water Equations
The final example that we consider is the two-dimensional nonlinear shallow water equations with three coupled fields: fluid column height $\eta(x,y,t)$ and velocity components $u(x,y,t)$ and $v(x,y,t)$. These equations are derived by depth-integrating the Navier-Stokes equations under the assumption that horizontal length scales greatly exceed vertical scales. This leads to nearly hydrostatic vertical pressure gradients and depth-constant horizontal velocities, allowing the vertical velocity to be eliminated through integration. While vertical velocity is removed from the equations, it can be recovered from the continuity equation once the horizontal solution is obtained. The governing equations are:
$\dfrac{\partial(\eta)}{\partial t}+\dfrac{\partial(\eta u)}{\partial x}+\dfrac{\partial(\eta v)}{\partial y}=0,$ (8)
$\dfrac{\partial(\eta u)}{\partial t}+\dfrac{\partial}{\partial x}\left(\eta u^{2}+\dfrac{1}{2}g\eta^{2}\right)+\dfrac{\partial(\eta uv)}{\partial y}=\nu(u_{xx}+u_{yy}),$ (9)
$\dfrac{\partial(\eta v)}{\partial t}+\dfrac{\partial(\eta uv)}{\partial x}+\dfrac{\partial}{\partial y}\left(\eta v^{2}+\dfrac{1}{2}g\eta^{2}\right)=\nu(v_{xx}+v_{yy}).$ (10)
The system is solved on the spatiotemporal domain defined as: $(x,y)\in[0,1]\times[0,1]$ and $t\in[0,1]$. The initial conditions are defined as follows. The total fluid column height $\eta(x,y,t)$, is given by a mean value of 1 plus some initial perturbation around a unit mean depth, $\eta_{0}(x,y)$. The initial velocity fields $u(x,y,t)$ and $v(x,y,t)$ are zero. We set the gravitational acceleration $g=1$ and the viscosity coefficient $\nu=0.002$. This problem presents a particularly challenging benchmark that combines multiple coupled fields with complex nonlinear dynamics. The shallow water equations serve as an excellent test case due to their physical relevance in modeling tsunami propagation and other geophysical flows. Further, these equations share structural similarities with compressible fluid dynamics, differing primarily in the absence of an energy equation dependent on the equation of state.
Mathematically, the shallow water equations form a nonlinear hyperbolic system of conservation laws. The advective fluxes capture nonlinear momentum transport, while the pressure-gradient and source terms govern gravity-driven wave propagation and topographic interactions. The dynamics are characterized by the coexistence of smooth wave-like solutions and sharp gradients, including hydraulic jumps, shocks, and bore formations. This multiscale behavior leads to complex flow patterns that are highly sensitive to initial and boundary conditions. For surrogate modeling, the shallow water equations pose two key challenges: (i) the strong nonlinearity of the advective terms amplifies errors over long-horizon rollouts, particularly in the presence of shocks and discontinuities, and (ii) the simultaneous need to capture fine-scale features such as bores and fronts, alongside basin-scale wave propagation and circulation, demands models that are robust across multiple spatial and temporal scales.
Figure 8: 2D nonlinear shallow water equations: Performance of all frameworks on 50 training trajectories, evaluated on 500 trajectories for predicting the $u$-velocity field of a representative sample. Top row: ground truth solution contours; subsequent rows: absolute error contours for each framework.
Figure 9: 2D nonlinear shallow water equations: Performance of all frameworks on 50 training trajectories, evaluated on 500 trajectories for predicting the $v$-velocity field of a representative sample. Top row: ground truth solution contours; subsequent rows: absolute error contours for each framework.
Figure 10: 2D nonlinear shallow water equations: Performance of all frameworks on 50 training trajectories, evaluated on 500 trajectories for predicting the height field, $h$, of a representative sample. Top row: ground truth solution contours; subsequent rows: absolute error contours for each framework.
The ground truth dataset is generated following a strategy similar to those in cases 3.1 and 3.2. As before, the initial conditions are sampled from smooth Gaussian random fields with Matérn-type kernels and periodic boundary conditions. Once again, to pick the most informative subset of training trajectories, the PCA-KMeans procedure was employed (see Section 3.1) by retaining the first 50 dominant principal vectors following the PCA dimensionality reduction step.
We now examine the performance of different surrogates for the three coupled fields: the velocity components $u$ and $v$, and the height field $h$. Figures 8, 9, and 10 present a qualitative comparison of prediction errors for GNS, FNO AR, and FNO FR. All methods were trained on 50 trajectories and evaluated on 500 test trajectories. Notably, the DeepONet variants exhibited poor performance on this problem, producing errors several orders of magnitude larger than the FNO variants and GNS. This degraded performance can be attributed to the initial conditions where both velocity fields $u$ and $v$ are exactly zero, creating a potentially challenging scenario for DeepONet’s operator learning framework. Hence, for brevity, we focus our analysis only on the comparison between GNS and the two FNO variants (FNO AR and FNO FR). Between the FNO variants, FNO FR incurs larger errors than FNO AR. This can be attributed to the inherent difficulty of the FNO FR framework in trying to learn the solution operator from near-zero initial conditions (exactly zero $u$ and $v$ velocity fields). GNS performs the best out of all three frameworks, with the errors in a very small range and the predictions showing a near perfect match, with a modest training set of 50 trajectories. The rationale is that in spite of the near zero initial conditions, the evolution dynamics of $u$ and $v$ velocity fields is effortlessly handled by GNS. Primarily in the SWE case, the initial perturbation provided to the height field is what sets up or induces the velocity fields and these nonlinear interactions are extremely well captured by GNS owing to its ability to model such interactions in a more physically-consistent manner that closely mimicks the true dynamics.
Table 5: Two-Dimensional nonlinear shallow water equations with vector output field: Inference errors for $u$, $v$, and $h$ fields across different methods and training trajectory counts. At inference, 500 trajectories are used, with a timestep of $\Delta t=0.01$ for GNS.
Method Training Trajectories
30 50 70 100
Field $u$
GNS 0.0323 0.01718 0.01175 0.00505
FNO AR 0.0347 0.03682 0.02904 0.02366
FNO FR 0.1496 0.09056 0.06069 0.04432
Field $v$
GNS 0.0270 0.01916 0.00933 0.00468
FNO AR 0.03562 0.03098 0.02373 0.01947
FNO FR 0.1343 0.08105 0.05415 0.03989
Field $h$
GNS 0.00074 0.00048 0.00021 0.00012
FNO AR 0.00102 0.00097 0.00076 0.00062
FNO FR 0.00411 0.00246 0.00162 0.00117
Table 5 presents a comprehensive evaluation of the relative $L_{2}$ test errors for GNS and FNO-based variants trained with subsets of 30, 50, 70, and 100 trajectories. The quantitative error values exhibit strong agreement with the qualitative observations in Figs. 8, 9, and 10. Across all training subsets, GNS consistently delivers the best performance, achieving the lowest errors for all three coupled fields, $u$, $v$, and $h$. This is closely followed by FNO AR, which has slightly higher errors than GNS but performs substantially better than FNO FR, which exhibits the highest errors across all training subsets. The primary reason for the deterioration in FNO FR performance is the underlying operator learning problem it faces: mapping near-zero initial conditions (due to initial zero velocities) to the full spatiotemporal solution field, which, as established earlier, is an especially challenging task.
Figure 11: 2D nonlinear shallow water equations: Temporal evolution of error accumulation across GNS (black), FNO AR (orange), and FNO FR (blue).
We also assess the temporal error growth for the GNS and FNO frameworks in Fig. 11. In addition to exhibiting higher generalization accuracy in the limited-data regime, GNS (black) demonstrates stable and controlled error growth, indicating robustness and reliability in predicting accurate solutions over extended time horizons. FNO AR (orange) follows closely, while FNO FR (blue) shows comparatively higher error accumulation, consistent with the results in Table 5. Finally, the relatively large $L_{2}$ errors for the $u$ and $v$ velocity fields in Fig. 11 for FNO FR can be attributed to the zero initial conditions and the difficulty FNO FR encounters in learning the coupling between the initial height perturbation and the subsequent evolution of the velocity fields. By contrast, GNS is able to effectively capture this coupling despite the zero initial velocity fields.